OpenStudy (anonymous):

What is the acceleration of a cabinet of mass 45 kilograms if Jake and Ted push it by applying horizontal force of 25 newtons and 18 newtons respectively in the same direction? (Assume a flat, frictionless surface.)

OpenStudy (anonymous):

f1+f2=ma 25+18=45a 43=45a a=43/45

OpenStudy (anonymous):

so wats the answer cause that what i need

OpenStudy (anonymous):

.

OpenStudy (stormfire1):

pastory posted the answer....what part don't you understand about it?

OpenStudy (anonymous):

a=43/45 a=0.956